Monthly Archives: August 2014

140821 LML

Robert arrived at LML at 9:15. The doors were open but no sample had been collected. The jar had disconnected from the collector. It was reattached and the top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.

140821 UCSC

Kona got to the roof od ISB at 9:30. No sample had been collected by the collector nor the new funnel apparatus. She washed the collector in place with soap and water, ran a successful test run, then activated the system for fog collection.

140820 UCSC Observations

Isaac arrived at ISB at 9:45am. Upon arrival there was a light drizzle coming down. The doors to the collector had been deployed and the big fag was on. However, upon examination of the collection jar no sample was present. The collector was then disassembled and its components as well as the collector body were washed with 409 and DI water. The system was then reset for the next fog collection event. In addition due to the presence of precipitation a passive rain collector was set up in order to collect any rainfall.

140817 LML Observations

Isaac arrived at LML at 9:30am. The doors to the collector had been deployed but no sample was present in the collection jar. The collector was then disassembled and the body of the collector as well as its components were sprayed initially with 409 followed by a rinse with DI water. The collector was then reassembled for the next fog event.

140816 LML Observations

Isaac arrived at LML at 10:30 am. The doors to the collector had been deployed but no sample was present in the collection jar. The collector was then disassembled and washed with 409 and DI water. Upon completion of this process the collector was reassembled for the next fog event.

140819ISB(UCSC)

Robert arrived at ISB at 10:15. The doors were open but no sample had been collected. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.

140819_Montara

Samples this morning and yesterday morning at Montara after two weeks of high and dry.

14/08/18: After one of our wetter nights lately, 50 mL sample found. Preserved and cooled on site. Removed all parts for thorough 409 wash, followed by copious water rinse and finally a DI rinse, followed by a DI spray. Attached new sample bottle.

14/08/19: A drippy morning, 30 mL sample, preserved and cooled on site, removed strand racks for thorough DI rinse throughout. Was still above 90% when I left so left it running.

140819 LML

Weird Santa Cruz fog.  Dry all night, then started drizzling both at the seashore and at 800′ elevation starting at about 8 am.  Ground noticeably wet in openings.  No samples at either locations.  Collector at LML seems to be in working order (program, latches, etc).  Rinsed with DI water with strands in place.  Rinsed and reused sample jar.

140818 R/V Point Sur

Positioned south of the outer Monterey canyon on the night of August 17th, 2014 a noticeably wet fog set in with light winds from the west.  We cleaned the collector, took an equipment blank and began collecting fog at 2315 hrs with ship maintaining position into the wind.  We checked sample at 0000 hrs on August 18th and noted about 10 mL of sample.  Continued collecting and checked sample at 0200 with little change in volume.   We checked sample at 0400 hrs and noted no change in sample volume.  Checked sample again at 0600 hrs and noted no change in sample.  At 0840 hrs we headed east toward our Brazil Ranch station and continued collecting sample while in transit.  At 1025 hrs we shut down the collector and retrieved a modest ~20 mls of fog water and making it a total of four fog samples collected on this cruise to date.